OpenStudy (anonymous):

Solve the system algebraically: 32x^2 + 9y^2 = 324 3y^2 - x^2 = 3

OpenStudy (blake57roger):

32x^2+9y^2=324 3y^2-x^2=3 this appears to be 2 equations ... Solve for the intersections? 3(3y^2-x^2=3 ) --- 9y^2 = 3x^2 +9 << sub into other equation 32x^2 + 3x^2 + 9 =324 35x^2 = 315 x^2 = 9 --- x = +/- 3 3y^2-x^2=3 for x = 3 --- 3y^2 = 12 --- y = +/-2 for x = -3 --- 3y^2 = 12 --- y = +/-2 (3,2), (3,-2), (-3,2), (-3,-2) << answers

OpenStudy (anonymous):

(3,2), (3,-2), (-3,2), (-3,-2)

OpenStudy (anonymous):

@blake57roger dont copy off the internet that's playgerisem

OpenStudy (blake57roger):

just the anwsers are the way you do it isnt

OpenStudy (blake57roger):

medal lol

OpenStudy (anonymous):

dont copy other people